# Structured Probabilistic Inference & Generative Modeling workshop of NeurIPS 2025

### v2

@@ D.3 Proof of Corollary 1
For reverse KL divergence, $f(t)=-\log(t)$, so $f^{\prime}(t)=-1/t$. Therefore:
$w_{f}(q_{y}(i),p(i|x))=p(i|x)\cdot f^{\prime}\left(\frac{q_{y}(i)}{p(i|x)}\right)=p(i|x)\cdot\left(-\frac{p(i|x)}{q_{y}(i)}\right)=-p(i|x)\cdot\frac{p(i|x)}{q_{y}(i)}.$ (62)
Wait, let me recalculate. For reverse KL:
$D_{\text{KL}}(q_{y}\|p)=\sum_{i}q_{y}(i)\log\frac{q_{y}(i)}{p(i|x)}=\sum_{i}q_{y}(i)\log q_{y}(i)-\sum_{i}q_{y}(i)\log p(i|x).$ (63)
Taking gradient with respect to $x$ (and noting $q_{y}$ is independent of $x$):
$\nabla_{x}D_{\text{KL}}(q_{y}\|p)=-\sum_{i}q_{y}(i)\nabla_{x}\log p(i|x)=-\sum_{i}q_{y}(i)\left[\nabla_{x}f_{i}(x)-\sum_{j}p(j|x)\nabla_{x}f_{j}(x)\right].$ (64)
This gives $w_{f}(q_{y}(i),p(i|x))=q_{y}(i)$ for reverse KL. Substituting into Equation (8) and noting the sign flip from $\alpha\nabla_{x}D_{f}$ yields the corollary. ∎
@@ D.4 Proof of Corollary 2
For forward KL divergence, $f(t)=t\log(t)$, so $f^{\prime}(t)=\log(t)+1$. The divergence is:
$D_{\text{KL}}(p\|q_{y})=\sum_{i}p(i|x)\log\frac{p(i|x)}{q_{y}(i)}.$ (65)
Taking the gradient:
$\nabla_{x}D_{\text{KL}}(p\|q_{y})$ $=\sum_{i}\nabla_{x}p(i|x)\log\frac{p(i|x)}{q_{y}(i)}+\sum_{i}p(i|x)\nabla_{x}\log\frac{p(i|x)}{q_{y}(i)}$ (66)
$=\sum_{i}p(i|x)\nabla_{x}\log p(i|x)\left[\log\frac{p(i|x)}{q_{y}(i)}+1\right]$ (67)
$=\sum_{i}p(i|x)\left[\nabla_{x}f_{i}(x)-\sum_{j}p(j|x)\nabla_{x}f_{j}(x)\right]\left[\log\frac{p(i|x)}{q_{y}(i)}+1\right].$ (68)
Rearranging and noting the sign for $\alpha\nabla_{x}D_{\text{KL}}(p\|q_{y})$ in our formulation yields the corollary. ∎
@@ D.5 Proof of Corollary 3
Jensen-Shannon divergence is defined as:
$D_{\text{JS}}(q_{y}\|p)=\frac{1}{2}D_{\text{KL}}(q_{y}\|m)+\frac{1}{2}D_{\text{KL}}(p\|m),$ (69)
where $m(i)=\frac{1}{2}(q_{y}(i)+p(i|x))$. Taking the gradient with respect to $x$:
$\nabla_{x}D_{\text{JS}}(q_{y}\|p)$ $=\frac{1}{2}\nabla_{x}D_{\text{KL}}(q_{y}\|m)+\frac{1}{2}\nabla_{x}D_{\text{KL}}(p\|m)$ (70)
$=-\frac{1}{2}\sum_{i}q_{y}(i)\nabla_{x}\log m(i)+\frac{1}{2}\sum_{i}\left[\nabla_{x}p(i|x)\log\frac{p(i|x)}{m(i)}+p(i|x)\nabla_{x}\log\frac{p(i|x)}{m(i)}\right].$ (71)
Since $\nabla_{x}\log m(i)=\frac{1}{2m(i)}\nabla_{x}p(i|x)=\frac{p(i|x)}{2m(i)}\nabla_{x}\log p(i|x)$ and using $\nabla_{x}\log\frac{p(i|x)}{m(i)}=\nabla_{x}\log p(i|x)-\nabla_{x}\log m(i)$, we can simplify:
$\nabla_{x}D_{\text{JS}}(q_{y}\|p)$ $=\frac{1}{2}\sum_{i}\left[\frac{p(i|x)-q_{y}(i)}{m(i)}\right]p(i|x)\nabla_{x}\log p(i|x)$ (72)
$=\frac{1}{2}\sum_{i}\left[\frac{p(i|x)-q_{y}(i)}{m(i)}\right]p(i|x)\left[\nabla_{x}f_{i}(x)-\sum_{j}p(j|x)\nabla_{x}f_{j}(x)\right].$ (73)
Including the sign from $\alpha\nabla_{x}D_{\text{JS}}$ yields the corollary. ∎
@@ D.6 Proof of Lemma 1
Direct consequence of the reverse KL gradient formula:
$\nabla_{x}D_{\text{KL}}(q_{y}\|p)=-\sum_{i}q_{y}(i)\nabla_{x}\log p(i|x)=-\sum_{i}q_{y}(i)\left[\nabla_{x}f_{i}(x)-\sum_{j}p(j|x)\nabla_{x}f_{j}(x)\right],$ (74)
which can be rewritten by distributing the sum:
$\sum_{i}q_{y}(i)\nabla_{x}f_{i}(x)+\sum_{i}q_{y}(i)\sum_{j}p(j|x)\nabla_{x}f_{j}(x)=-\sum_{i}q_{y}(i)\nabla_{x}f_{i}(x)+\sum_{j}p(j|x)\nabla_{x}f_{j}(x),$ (75)
where we used $\sum_{i}q_{y}(i)=1$. Flipping the sign gives the decomposition. ∎
@@ D.7 Proof of Proposition 3
For Gaussian components with $\nabla_{x}f_{k}(x)=\Sigma_{k}^{-1}(\mu_{k}-x)$ and $p(k|x)=w_{k}(x)$, substituting into Corollary 1:
$\nabla_{x}\mathcal{S}_{\text{RKL}}(x,y)$ $=\tau_{1}\Sigma_{y}^{-1}(\mu_{y}-x)-\tau_{2}\sum_{k=1}^{K}p_{\tau_{2}}(k|x)\Sigma_{k}^{-1}(\mu_{k}-x)$ (76)
$\quad+\alpha\sum_{k=1}^{K}q_{y}(k)\left[\Sigma_{k}^{-1}(\mu_{k}-x)-\sum_{j=1}^{K}w_{j}(x)\Sigma_{j}^{-1}(\mu_{j}-x)\right].$ (77)
Factoring $\Sigma_{k}^{-1}(\mu_{k}-x)$ from each term and collecting coefficients yields the weight function $\Gamma_{k}(x,y,\alpha,\epsilon)$ as stated. ∎
@@ D.8 Proof of Corollary 4
For squared Hellinger distance, we use the definition:
$D_{H^{2}}(q_{y}\|p(\cdot|x))=\sum_{i=1}^{N}\left(\sqrt{q_{y}(i)}-\sqrt{p(i|x)}\right)^{2}.$ (78)
Taking the gradient with respect to $x$ (noting that $q_{y}$ is independent of $x$):
$\nabla_{x}D_{H^{2}}(q_{y}\|p(\cdot|x))$ $=\sum_{i=1}^{N}2\left(\sqrt{q_{y}(i)}-\sqrt{p(i|x)}\right)\cdot\nabla_{x}\left(-\sqrt{p(i|x)}\right)$ (79)
$=\sum_{i=1}^{N}2\left(\sqrt{q_{y}(i)}-\sqrt{p(i|x)}\right)\cdot\left(-\frac{1}{2\sqrt{p(i|x)}}\right)\nabla_{x}p(i|x)$ (80)
$=-\sum_{i=1}^{N}\frac{\sqrt{q_{y}(i)}-\sqrt{p(i|x)}}{\sqrt{p(i|x)}}p(i|x)\nabla_{x}\log p(i|x)$ (81)
$=-\sum_{i=1}^{N}\left(\sqrt{q_{y}(i)p(i|x)}-p(i|x)\right)\nabla_{x}\log p(i|x).$ (82)
Substituting $\nabla_{x}\log p(i|x)=\nabla_{x}f_{i}(x)-\sum_{j}p(j|x)\nabla_{x}f_{j}(x)$:
$\nabla_{x}D_{H^{2}}(q_{y}\|p(\cdot|x))$ $=-\sum_{i=1}^{N}\left(\sqrt{q_{y}(i)p(i|x)}-p(i|x)\right)\left[\nabla_{x}f_{i}(x)-\sum_{j=1}^{N}p(j|x)\nabla_{x}f_{j}(x)\right].$ (83)
Observe that the term involving $p(i|x)$ vanishes:
$\sum_{i=1}^{N}p(i|x)\left[\nabla_{x}f_{i}(x)-\sum_{j=1}^{N}p(j|x)\nabla_{x}f_{j}(x)\right]=\sum_{i=1}^{N}p(i|x)\nabla_{x}f_{i}(x)-\sum_{j=1}^{N}p(j|x)\nabla_{x}f_{j}(x)=0.$ (84)
Therefore:
$\nabla_{x}D_{H^{2}}(q_{y}\|p(\cdot|x))=-\sum_{i=1}^{N}\sqrt{q_{y}(i)p(i|x)}\left[\nabla_{x}f_{i}(x)-\sum_{j=1}^{N}p(j|x)\nabla_{x}f_{j}(x)\right].$ (85)
Including the negative sign from $\alpha\nabla_{x}D_{H^{2}}$ in our score formulation $\mathcal{S}_{H^{2}}(x,y)=\log p_{\tau_{1},\tau_{2}}(y|x)-\alpha D_{H^{2}}(q_{y}\|p(\cdot|x))$ and combining with the classifier guidance term yields:
$\quad+\alpha\sum_{i=1}^{N}\sqrt{q_{y}(i)p(i|x)}\left[\nabla_{x}f_{i}(x)-\sum_{j=1}^{N}p(j|x)\nabla_{x}f_{j}(x)\right],$
which can be rewritten in the form stated in the corollary by factoring out the square root terms. ∎

### v3

@@ D.3 Proof of Corollary 1
Recall Proposition 2 gives
$\nabla_{x}\mathcal{S}(x,y)=\tau_{1}\nabla_{x}f_{y}(x)-\tau_{2}\sum_{i=1}^{N}p_{\tau_{2}}(i|x)\,\nabla_{x}f_{i}(x)-\alpha\nabla_{x}D_{f}(q_{y}\|p(\cdot|x)),$ (57)
and for an $f$-divergence of the form $D_{f}(q\|p)=\sum_{i}q(i)f\!\left(\frac{p(i)}{q(i)}\right)$,
$\nabla_{x}D_{f}(q_{y}\|p(\cdot|x))=\sum_{i=1}^{N}w_{f}\big(q_{y}(i),p(i|x)\big)\,g_{i}(x),\qquad w_{f}(q,p)=p\,f^{\prime}\!\left(\frac{p}{q}\right),$ (58)
where $g_{i}(x)=\nabla_{x}\log p(i|x)$.
For reverse KL, $D_{\mathrm{KL}}(q_{y}\|p)=\sum_{i}q_{y}(i)\log\frac{q_{y}(i)}{p(i|x)}$ corresponds to $f(t)=-\log t$ with $t=\frac{p}{q}$. Hence $f^{\prime}(t)=-1/t$ and
$w_{f}\big(q_{y}(i),p(i|x)\big)=p(i|x)\,f^{\prime}\!\left(\frac{p(i|x)}{q_{y}(i)}\right)=p(i|x)\left(-\frac{q_{y}(i)}{p(i|x)}\right)=-\,q_{y}(i).$ (59)
Therefore,
$\nabla_{x}D_{\mathrm{KL}}(q_{y}\|p(\cdot|x))=\sum_{i=1}^{N}\big(-q_{y}(i)\big)\,g_{i}(x)=-\sum_{i=1}^{N}q_{y}(i)\,g_{i}(x),$ (60)
and the regularizer contribution in the score is $\alpha\nabla_{x}D_{\mathrm{KL}}(q_{y}\|p)=+\alpha\sum_{i}q_{y}(i)g_{i}(x).$ Substituting this into Proposition 2 yields Corollary 1. ∎
@@ D.4 Proof of Corollary 2
Forward KL is $D_{\mathrm{KL}}(p\|q_{y})=\sum_{i}p(i|x)\log\frac{p(i|x)}{q_{y}(i)}$. With our $f$-divergence convention $D_{f}(q\|p)=\sum_{i}q(i)f\!\left(\frac{p(i)}{q(i)}\right)$, this corresponds to $f(t)=t\log t$ (again with $t=\frac{p}{q}$). Indeed, $\sum_{i}q(i)\,t\log t=\sum_{i}p(i)\log\frac{p(i)}{q(i)}.$ Then $f^{\prime}(t)=\log t+1$, and the associated weight is
$w_{f}\big(q_{y}(i),p(i|x)\big)=p(i|x)\Big(\log\frac{p(i|x)}{q_{y}(i)}+1\Big).$ (61)
Hence,
$\nabla_{x}D_{\mathrm{KL}}(p\|q_{y})=\sum_{i=1}^{N}p(i|x)\Big(\log\frac{p(i|x)}{q_{y}(i)}+1\Big)\,g_{i}(x),\qquad g_{i}(x)=\nabla_{x}\log p(i|x),$ (62)
and plugging into Proposition 2 gives the regularizer term $\alpha\nabla_{x}D_{\mathrm{KL}}(p\|q_{y})=-\alpha\sum_{i}p(i|x)\big(\log\frac{p(i|x)}{q_{y}(i)}+1\big)g_{i}(x),$ which is exactly Corollary 2. ∎
@@ D.5 Proof of Corollary 3
The Jensen–Shannon divergence is
$D_{\mathrm{JS}}(q_{y}\|p)=\frac{1}{2}D_{\mathrm{KL}}(q_{y}\|m)+\frac{1}{2}D_{\mathrm{KL}}(p\|m),\qquad m(i)=\frac{q_{y}(i)+p(i|x)}{2}.$ (63)
Since $q_{y}$ is fixed, only $p(\cdot|x)$ (and thus $m$) depends on $x$. Differentiate $D_{\mathrm{JS}}$ w.r.t. $p(i|x)$ (treating $p$ as free variables first). A direct computation gives the partial derivative
$\frac{\partial}{\partial p(i)}D_{\mathrm{JS}}(q_{y}\|p)=\frac{1}{2}\log\frac{p(i)}{m(i)}.$ (64)
(Indeed, the $q_{y}\|m$ term contributes $\frac{q_{y}(i)}{4m(i)}$, the $p\|m$ term contributes $\frac{1}{2}\big(\log\frac{p(i)}{m(i)}+1\big)-\frac{p(i)}{4m(i)}$, and the rational terms cancel because $q_{y}(i)+p(i)=2m(i)$.)
Therefore, by the chain rule,
$\nabla_{x}D_{\mathrm{JS}}(q_{y}\|p(\cdot|x))=\sum_{i=1}^{N}\frac{1}{2}\log\frac{p(i|x)}{m(i)}\,\nabla_{x}p(i|x)=\sum_{i=1}^{N}\frac{p(i|x)}{2}\log\frac{p(i|x)}{m(i)}\,g_{i}(x),$ (65)
where $\nabla_{x}p(i|x)=p(i|x)g_{i}(x)$ and $g_{i}(x)=\nabla_{x}\log p(i|x)$. Since $\frac{p(i|x)}{m(i)}=\frac{2p(i|x)}{p(i|x)+q_{y}(i)}$, this can be written as
$\nabla_{x}D_{\mathrm{JS}}(q_{y}\|p(\cdot|x))=\sum_{i=1}^{N}\frac{p(i|x)}{2}\log\frac{2p(i|x)}{p(i|x)+q_{y}(i)}\,g_{i}(x).$ (66)
Substituting into Proposition 2 and applying the $\alpha$ factor in the score yields Corollary 3. ∎
@@ D.6 Proof of Lemma 1
By definition,
$D_{\mathrm{KL}}(q_{y}\|p(\cdot|x))=\sum_{i=1}^{N}q_{y}(i)\log\frac{q_{y}(i)}{p(i|x)}=\text{const}-\sum_{i=1}^{N}q_{y}(i)\log p(i|x),$ (67)
so
$\nabla_{x}D_{\mathrm{KL}}(q_{y}\|p(\cdot|x))=\sum_{i=1}^{N}q_{y}(i)\,\nabla_{x}\log p(i|x).$ (68)
For softmax probabilities $p(i|x)=\frac{e^{f_{i}(x)}}{\sum_{k}e^{f_{k}(x)}}$,
$\nabla_{x}\log p(i|x)=\nabla_{x}f_{i}(x)-\sum_{j=1}^{N}p(j|x)\nabla_{x}f_{j}(x).$ (69)
Multiplying by $q_{y}(i)$ and summing over $i$ gives
$-\nabla_{x}D_{\mathrm{KL}}(q_{y}\|p(\cdot|x))$ $=\sum_{i=1}^{N}q_{y}(i)\nabla_{x}f_{i}(x)-\Big(\sum_{i=1}^{N}q_{y}(i)\Big)\sum_{j=1}^{N}p(j|x)\nabla_{x}f_{j}(x)$
$=\sum_{i=1}^{N}q_{y}(i)\nabla_{x}f_{i}(x)-\sum_{j=1}^{N}p(j|x)\nabla_{x}f_{j}(x),$ (70)
since $\sum_{i}q_{y}(i)=1$. This is the claimed decomposition. ∎
@@ D.7 Proof of Proposition 3
Let $\ell_{k}(x)=\log b_{k}+\log f_{k}(x)$. For Gaussian components, $\nabla_{x}\ell_{k}(x)=\nabla_{x}\log f_{k}(x)=\Sigma_{k}^{-1}(\mu_{k}-x)$. Additionally, since $\nabla_{x}\log f_{k}(x)=\frac{1}{f_{k}(x)}\nabla_{x}f_{k}(x)\rightarrow\nabla_{x}f_{k}(x)=f_{k}(x)\nabla_{x}\log f_{k}(x)$. Substituting these into Corollary 1 gives
$\nabla_{x}\mathcal{S}_{\text{RKL}}(x,y)$ $=\tau_{1}f_{y}(x)\nabla_{x}\ell_{y}(x)-\tau_{2}\sum_{k=1}^{K}p_{\tau_{2}}(k|x)f_{k}(x)\nabla_{x}\ell_{k}(x)$ (71)
$\quad+\alpha\sum_{k=1}^{K}q_{y}(k)\Big(f_{k}(x)\nabla_{x}\ell_{k}(x)-\sum_{j=1}^{K}f_{j}(x)w_{j}(x)\nabla_{x}\ell_{j}(x)\Big)$
$=\sum_{k=1}^{K}f_{k}(x)\Sigma_{k}^{-1}(\mu_{k}-x)$
$\times\Big[\tau_{1}\mathbb{I}_{k=y}-\tau_{2}p_{\tau_{2}}(k|x)+\alpha q_{y}(k)\Big]$
$\;-\;\alpha\Big(\sum_{k=1}^{K}q_{y}(k)\Big)\sum_{j=1}^{K}f_{j}(x)w_{j}(x)\Sigma_{j}^{-1}(\mu_{j}-x).$
Since $\sum_{k}q_{y}(k)=1$, collecting the coefficient of each $f_{k}(x)\Sigma_{k}^{-1}(\mu_{k}-x)$ yields $\Gamma_{k}(x,y,\alpha,\epsilon)=\tau_{1}\mathbb{I}_{k=y}-\tau_{2}p_{\tau_{2}}(k|x)+\alpha(q_{y}(k)-w_{k}(x))$ as stated. ∎
@@ D.8 Proof of Corollary 4
The squared Hellinger distance can be written as
$D_{H^{2}}(q_{y}\|p(\cdot|x))=\sum_{i=1}^{N}\big(\sqrt{q_{y}(i)}-\sqrt{p(i|x)}\big)^{2}=\sum_{i}q_{y}(i)+\sum_{i}p(i|x)-2\sum_{i}\sqrt{q_{y}(i)p(i|x)}.$ (72)
Since $\sum_{i}q_{y}(i)=1$ and $\sum_{i}p(i|x)=1$, this reduces to $D_{H^{2}}(q_{y}\|p)=2-2\sum_{i}\sqrt{q_{y}(i)p(i|x)}.$ Thus,
$\nabla_{x}D_{H^{2}}(q_{y}\|p(\cdot|x))=-2\sum_{i=1}^{N}\nabla_{x}\sqrt{q_{y}(i)p(i|x)}=-2\sum_{i=1}^{N}\frac{\sqrt{q_{y}(i)}}{2\sqrt{p(i|x)}}\,\nabla_{x}p(i|x).$ (73)
Using $\nabla_{x}p(i|x)=p(i|x)\,g_{i}(x)$ with $g_{i}(x)=\nabla_{x}\log p(i|x)$,
$\nabla_{x}D_{H^{2}}(q_{y}\|p(\cdot|x))=-\sum_{i=1}^{N}\sqrt{q_{y}(i)p(i|x)}\,g_{i}(x).$ (74)
Finally, for the score $\mathcal{S}_{H^{2}}(x,y)=\log p_{\tau_{1},\tau_{2}}(y|x)-\alpha D_{H^{2}}(q_{y}\|p(\cdot|x)),$ we obtain
$\nabla_{x}\mathcal{S}_{H^{2}}(x,y)=\tau_{1}\nabla_{x}f_{y}(x)-\tau_{2}\sum_{i=1}^{N}p_{\tau_{2}}(i|x)\nabla_{x}f_{i}(x)+\alpha\sum_{i=1}^{N}\sqrt{q_{y}(i)p(i|x)}\,g_{i}(x),$ (75)
which is the stated corollary (after rewriting $g_{i}(x)$ via the softmax identity if desired). ∎